# Radiative corrections to superallowed $\boldsymbol{\beta}$ decays in effective field theory

### v1

@@ Abstract
Based on the EFT framework described here, we have derived a master formula for the decay rate and performed first numerical calculations for $\delta_{\text{NS}}$ in the decay ${}^{14}\text{O}\to{}^{14}\text{N}$ with Quantum Monte Carlo methods, confirming the expectations from the EFT power counting Cirigliano et al. . As an illustration, we extract $V_{ud}$ from the ${}^{14}\text{O}$ decay, finding $V_{ud}\big{[}{}^{14}\text{O}\big{]}=0.97364(56)$, with uncertainty dominated by our ignorance of the LECs, $(\delta V_{ud})_{g_{V}^{N\!N}}=4.3\times 10^{-4}$. Eliminating this uncertainty would result in $\delta V_{ud}=3.6\times 10^{-4}$. This is to be compared with $V_{ud}\big{[}{}^{14}\text{O}\big{]}=0.97405(37)$ from Ref. Hardy and Towner (2020), with uncertainty dominated by $\delta_{\text{NS}}$, $(\delta V_{ud})_{\delta_{\text{NS}}}=3.1\times 10^{-4}$, and with $V_{ud}=0.97373(31)$ obtained by a global analysis of the $0^{+}\to 0^{+}$ decays Hardy and Towner (2020). These considerations show that there is a clear path towards reaching $\delta V_{ud}\simeq 3\times 10^{-4}$, once the LECs are determined following the strategies outlined above. We expect that a few decays of light nuclei, combined with nuclear-structure calculations, should suffice to obtain a competitive determination of $V_{ud}$, including a robust estimate of the nuclear-structure uncertainties.
Discussion and outlook.—We have performed a first study of RC to superallowed nuclear $\beta$ decays in an EFT framework that bridges the EW scale to nuclear scales. We have identified the leading nuclear-structure-dependent corrections $\delta_{\text{NS}}$ as arising from matrix elements of EW transition operators of $\mathcal{O}(G_{F}\alpha\epsilon_{\not{\pi}},G_{F}\alpha\epsilon_{\chi})$ between initial and final nuclear wave functions. Several terms, such as the magnetic and recoil pieces of $\delta_{\text{NS}}$, already appear in the seminal work Towner (1992), while others are new. Most strikingly, we identified novel pion-exchange and short-range corrections that affect $\delta_{\text{NS}}$ at the same order as the usually considered corrections.

### v2

@@ Abstract
Based on the EFT framework described here, we have derived a master formula for the decay rate and performed first numerical calculations for $\delta_{\text{NS}}$ in the decay ${}^{14}\text{O}\to{}^{14}\text{N}$ with Quantum Monte Carlo methods, confirming the expectations from the EFT power counting Cirigliano et al. (2024b). As an illustration, we extract $V_{ud}$ from the ${}^{14}\text{O}$ decay, finding $V_{ud}\big{[}{}^{14}\text{O}\big{]}=0.97364(56)$, with uncertainty dominated by our ignorance of the LECs, $(\delta V_{ud})_{g_{V}^{N\!N}}=4.3\times 10^{-4}$. Eliminating this uncertainty would result in $\delta V_{ud}=3.6\times 10^{-4}$. This is to be compared with $V_{ud}\big{[}{}^{14}\text{O}\big{]}=0.97405(37)$ from Ref. Hardy and Towner (2020), with uncertainty dominated by $\delta_{\text{NS}}$, $(\delta V_{ud})_{\delta_{\text{NS}}}=3.1\times 10^{-4}$, and with $V_{ud}=0.97373(31)$ obtained by a global analysis of the $0^{+}\to 0^{+}$ decays Hardy and Towner (2020). These considerations show that there is a clear path towards reaching $\delta V_{ud}\simeq 3\times 10^{-4}$, once the LECs are determined following the strategies outlined above. We expect that a few decays of light nuclei, combined with nuclear-structure calculations, should suffice to obtain a competitive determination of $V_{ud}$, including a robust estimate of the nuclear-structure uncertainties.
Discussion and outlook.—We have performed a first study of RC to superallowed nuclear $\beta$ decays in an EFT framework that bridges the EW scale to nuclear scales. We have identified the leading nuclear-structure-dependent corrections $\delta_{\text{NS}}$ as arising from matrix elements of EW transition operators of $\mathcal{O}(G_{F}\alpha\epsilon_{\not{\pi}},G_{F}\alpha\epsilon_{\chi})$ between initial and final nuclear wave functions. Several terms, such as the magnetic and recoil pieces of $\delta_{\text{NS}}$, already appear in the seminal work Towner (1992), while others are new. Most strikingly, we identified novel pion-exchange and short-range corrections that affect $\delta_{\text{NS}}$ at the same order as the usually considered corrections. Furthermore, we have sketched a strategy using global fits to superallowed $\beta$ decays to empirically determine the contact operators’ Wilson coefficients.